# Normalized topological indices discriminate between architectures of branched macromolecules

### v1

@@ Results
Figure 4: Two methods of coarse-graining trees and their effects on branching topology. (a) $(\mathrm{ABC}^{*},H_{2}^{*})$ space of $10^{3}$ random trees and $10^{3}$ random Barabási-Albert (BA) trees with $N=100$ and the phase space of these trees after undergoing (b) modularity-maximization (MM) coarse-graining and (c) Laplacian renormalization group (LRG) coarse-graining reducing their size to $\widetilde{N}=30$. Example tree structures below panel (a) and their coarse-grained versions below panels (b) and (c) correspond to the points highlighted by black circles in the phase space. (d) Phase space flow of $10^{3}$ random trees upon their coarse-graining from $N=100$ to progressively smaller target sizes $\widetilde{N}$ using either MM (top row) or LRG (bottom row) coarse-graining. (e) Changes in the average values of normalized topological indices $\langle\mathrm{ABC}^{*}\rangle$ and $\langle H_{2}^{*}\rangle$ with the size of coarse-grained trees $\widetilde{N}$ using MM and LRG methods. The values of the two indices are averaged over $10^{3}$ random trees.
@@
Any coarse-graining by definition leads to a loss of information, and we can use the $(\mathrm{ABC}^{*},H_{2}^{*})$ space to examine in detail the consequences of coarse-graining for the reduced topology of branched macromolecules. For simplicity, we will start already from mathematical trees and explore the effect of different strategies to reduce their size and coarse-grain them into smaller trees. We focus on two examples of common graph coarse-graining methods, modularity-maximization (MM) coarse-graining and Laplacian renormalization group (LRG) coarse-graining (see Methods). Starting with trees of size $N=100$, we coarse-grain them to different target tree sizes $\widetilde{N}$ and study how their position in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space changes (Fig. 4). We do this both for general (random) tree topologies as well as for scale-free Barabási-Albert (BA) trees.
@@
First, we see that scale-free BA trees occupy only a very small part of the $(\mathrm{ABC}^{*},H_{2}^{*})$ space with high $H_{2}^{*}$ and $\mathrm{ABC}^{*}$ (Fig. 4a), which we have previously identified as corresponding to very star-like trees (Fig. 2e). Random trees are, on the other hand, distributed across the entire phase space of normalized indices, as we have already seen in Fig. 2. When both types of trees with initial size $N=100$ are coarse-grained to $30\%$ of their size (i.e., $\widetilde{N}=30$), the two coarse-graining methods produce quite different results. MM retains the approximate clustering of BA trees in the upper right corner of the phase space while compressing random trees to the left side of the space (Fig. 4b). LRG is less effective in preserving the topological properties as captured by the phase space (Fig. 4c): while they preserve the values of $H_{2}^{*}$ of BA trees, their values of $\mathrm{ABC}^{*}$ get spread out; random trees are squeezed to very low values of $\mathrm{ABC}^{*}$. Example trees in Fig. 4 provide a visual intuition of the differences between the two methods: On the one hand, MM better preserves the global tree topology because it is optimised in terms of connectivity within a cluster. On the other hand, LRG is better at conserving the fine details of the branching topology because it optimises the clusters in terms of “diffusive” distance.
@@
We next progressively coarse-grain random trees towards smaller sizes $\widetilde{N}$ and monitor their flow in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space (panels (d) and (e) of Fig. 4). Again, MM better preserves the properties of random trees in the phase space compared to LRG, which always pushes trees towards small values of $\mathrm{ABC}^{*}$—i.e., towards more linear trees. Interestingly, the values of $\mathrm{ABC}^{*}$ of coarse-grained trees change non-monotonically with tree size when MM method is applied (Fig. 4e). When we monitor the same flow while coarse-graining scale-free BA trees (Fig. S10 in SI), their position in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space remains preserved well until the coarse-grained tree size $\widetilde{N}$ becomes very small; the only exception is again the effect of LRG on $\mathrm{ABC}^{*}$, as it decreases the its values quickly.
@@
Any coarse-graining applied to the structure of a branched macromolecule will necessarily change its underlying topology. Our framework—using the $(\mathrm{ABC}^{*},H_{2}^{*})$ space or even another combination of indices of interest to monitor the effects of coarse-graining methods on branching topology—provides a clear way to choose both the best coarse-graining strategy as well as the extent of coarse-graining most suited to the system and properties under examination.
Discussion
@@
In this work, we introduced a novel normalization of topological indices using the estimates of their cumulative distribution functions to create a two-dimensional phase space based on atom bond connectivity index $\mathrm{ABC}^{*}$ and second-order network coherence $H_{2}^{*}$. This phase space enables discrimination between branching architectures of macromolecules of different types and sizes. We demonstrated the usefulness of this phase space by exploring the topologies of different RNA structures mapped to trees using two different methods and by comparing the effects of two distinct tree coarse-graining methods on the resulting coarse-grained topologies.
@@
We have shown that the phase space of the two normalized topological indices $(\mathrm{ABC}^{*},H_{2}^{*})$ allows for a robust comparison of molecules of different sizes—stemming either from differences in their molecular weight or their mapping onto mathematical trees. Our choice of topological indices is important, as not all combinations of topological indices lead to such a well-defined and interpretable phase space. Figure S13 in SI shows an example of another phase space, defined through the commonly-used Wiener index $W$ and Randić index $R$, $(W^{*},R^{*})$ (see SI for their definitions). In this case, the normalization using Eq. (5) leads to a far less homogeneous phase space compared to $(\mathrm{ABC}^{*},H_{2}^{*})$ space. Approaches using a linear combination of several indices, extracted via principal component analysis, are also commonly used in QSPR studies [68, 69]; these, however, significantly impact both the interpretability of the phase space as well as the determination of the probability density functions needed for normalization.
@@
Methods
@@
The distribution parameters $p$ are first estimated with maximum likelihood estimation on sets of $10^{4}$ random trees at various $N$ ranging from $N=10$ to $N=300$ (Figs. S3 and S4 in SI). Afterwards, we fit the dependency of each distribution parameter $p$ on tree size $N$ with a power-law model of the form $p=\alpha N^{\beta}$ (Fig. S3). The power-law estimates of the distribution parameters are then used to obtain the cumulative distribution function of a given topological index for a tree of arbitrary size $N$.
@@
Conflict of interest
The authors declare no competing interest.
Data and code availability
Acknowledgements.
D.V. and A.B. acknowledge support by Slovenian Research Agency (ARRS) under contract no. P1-0055. J.M. and L.T. acknowledge support by ICSC—Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union—NextGenerationEU. J.M. acknowledges support by the European Union under NextGenerationEU grant PRIN 2022 PNRR Prot P2022TX4FE. J.M. and L.T. also thank the University of Trento for the strategic project “AIACE” that enabled this collaboration.

### v2

@@ IV Example: Coarse-graining the topologies of random trees
Figure 5: Two methods of coarse-graining trees and their effects on branching topology. (a) $(\mathrm{ABC}^{*},H_{2}^{*})$ space of $10^{3}$ random trees and $10^{3}$ random BA trees with $N=100$ and the phase space of these trees after undergoing (b) MM coarse-graining and (c) LRG coarse-graining reducing their size to $\widetilde{N}=30$. Example tree structures below panel (a) and their coarse-grained versions below panels (b) and (c) correspond to the points highlighted by black circles in the phase space. (d) Phase space flow of $10^{3}$ random trees upon their coarse-graining from $N=100$ to progressively smaller target sizes $\widetilde{N}$ using either MM (top row) or LRG (bottom row). (e) Changes in the average values of normalized topological indices $\langle\mathrm{ABC}^{*}\rangle$ and $\langle H_{2}^{*}\rangle$ and (f) changes in the average number of nodes with degree $i$, $\langle D_{i}\rangle$, with the size of coarse-grained trees $\widetilde{N}$ using MM and LRG. All averages are performed over $10^{3}$ random trees. Shaded regions indicate the range of one standard deviation from the mean.
@@
Any coarse-graining by definition leads to a loss of information, and we can use the $(\mathrm{ABC}^{*},H_{2}^{*})$ space to examine in detail the consequences of coarse-graining for the reduced topology of branched macromolecules. For simplicity, we will start already from mathematical trees and explore the effect of different strategies to reduce their size and coarse-grain them into smaller trees. We focus on two examples of common graph coarse-graining methods, modularity-maximization (MM) coarse-graining and Laplacian renormalization group (LRG) coarse-graining (see Appendix). Starting with trees of size $N=100$, we coarse-grain them to different target tree sizes $\widetilde{N}$ and study how their position in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space changes (Fig. 5). We do this both for general (random) tree topologies as well as for scale-free Barabási-Albert (BA) trees.
@@
First, we see that scale-free BA trees occupy only a very small part of the $(\mathrm{ABC}^{*},H_{2}^{*})$ space with high $H_{2}^{*}$ and $\mathrm{ABC}^{*}$ (Fig. 5a), which we have previously identified as corresponding to very star-like trees (Fig. 2e). Random trees are, on the other hand, distributed across the entire phase space of normalized indices, as we have already seen in Fig. 2. When both types of trees with initial size $N=100$ are coarse-grained to $30\%$ of their size (i.e., $\widetilde{N}=30$), the two coarse-graining methods produce quite different results. MM retains the approximate clustering of BA trees in the upper right corner of the phase space while compressing random trees to the left side of the space (Fig. 5b). LRG is less effective in preserving the topological properties as captured by the phase space (Fig. 5c): while they preserve the values of $H_{2}^{*}$ of BA trees, their values of $\mathrm{ABC}^{*}$ get spread out; random trees are squeezed to very low values of $\mathrm{ABC}^{*}$. Example trees in Fig. 5 provide a visual intuition of the differences between the two methods: On the one hand, MM better preserves the global tree topology because it is optimised in terms of connectivity within a cluster. On the other hand, LRG is better at conserving the fine details of the branching topology because it optimizes the clusters in terms of “diffusive” distance.
@@
We next progressively coarse-grain random trees towards smaller sizes $\widetilde{N}$ and monitor their flow in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space (panels (d) and (e) of Fig. 5). Again, MM better preserves the properties of random trees in the phase space compared to LRG, which always pushes trees towards small values of $\mathrm{ABC}^{*}$—i.e., towards more linear trees. This behaviour is even more obvious if we examine trajectories of individual trees during the coarse-graining process. Figure S9 in SI shows examples of four trees chosen from Fig. 5d, and we observe that LRG reduces the $\mathrm{ABC}^{*}$ value of trees to almost zero already at $80\%$ coarse-graining. This is not the case for MM, where the $\mathrm{ABC}^{*}$ values typically first increase and only gradually start moving towards smaller values as the trees are coarse-grained further. While $\mathrm{ABC}^{*}$ changes non-monotonically with MM coarse-graining, $H_{2}^{*}$ exhibits a plateau at coarse-graining levels below $\lesssim 70\%$ (Fig. 5e). When we monitor the same flow while coarse-graining scale-free BA trees (Fig. S10 in SI), their position in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space is preserved well until the coarse-grained tree size $\widetilde{N}$ becomes very small; the only exception is again the effect of LRG on $\mathrm{ABC}^{*}$, as the latter again decreases quickly. The different effects of MM and LRG on the topological indices of trees—and in particularly on $\mathrm{ABC}^{*}$—can be explained by taking a look into the distributions of node degrees in the coarse-grained trees (Fig. 5f). LRG initially increases the numbers of degree 2 and 3 nodes, and as we have already seen with RNA in Sec. III, a high proportion of degree $2$ nodes leads to very low values of $\mathrm{ABC}^{*}$. In contrast, MM at no point during the coarse-graining increases the number of nodes of any degree, but appears instead to initially preserve the number of high degree nodes in the coarse-grained trees.
@@
Any coarse-graining applied to the structure of a branched macromolecule will necessarily change its underlying topology. Our framework—using the $(\mathrm{ABC}^{*},H_{2}^{*})$ space or even another combination of indices of interest to monitor the effects of coarse-graining methods on branching topology—provides a clear way to choose both the best coarse-graining strategy as well as the extent of coarse-graining most suited to the system and properties under examination. While the choice of the best coarse-graining method strongly depends on its purpose, if the coarse-graining is meant to capture the essential features of tree topologies and retain the differences between them, a good method should maintain the position of the structures in the $(\mathrm{ABC}^{*},H_{2}^{*})$ space as much as possible during coarse-graining. Even though this calls for an in-depth study of different coarse-graining methods, our results (Fig. 5) indicate that between MM and LRG, this criterion strongly favours MM, with coarse-graining to $\sim 50\%$ of the initial tree size yielding the most consistent results.
V Discussion and conclusions
@@
In this work, we introduced a novel normalization of topological indices using the estimates of their cumulative distribution functions. We used this normalization to create a two-dimensional phase space based on atom bond connectivity index $\mathrm{ABC}^{*}$ and second-order network coherence $H_{2}^{*}$, which enables discrimination between branching architectures of macromolecules of different types and sizes. We demonstrated the usefulness of this phase space by exploring the topologies of different RNA structures mapped to trees using two different methods and by comparing the effects of two distinct tree coarse-graining methods on the resulting coarse-grained topologies.
@@
We have shown that the phase space of the two normalized topological indices $(\mathrm{ABC}^{*},H_{2}^{*})$ allows for a robust comparison of molecules of different sizes—stemming either from differences in their molecular weight or their mapping onto mathematical trees. Our choice of topological indices is important, as not all combinations of topological indices lead to such a well-defined and interpretable phase space. Figure S13 in SI shows an example of another phase space, defined through the commonly-used Wiener index $W$ and Randić index $R$, $(W^{*},R^{*})$ (see SI for their definitions). In this case, the normalization using Eq. (5) leads to a far less homogeneous phase space compared to $(\mathrm{ABC}^{*},H_{2}^{*})$ space, which comes in addition to the reduced discrimination power of these two indices (Fig. S6). Approaches using a linear combination of several indices, extracted via principal component analysis (PCA), are also commonly used in QSPR studies [68, 69]. Such approaches, however, significantly impact both the interpretability of the phase space spanned by the principal components as well as the determination of the probability density functions needed for normalization. Similarity between graph-like objects can also be studied using machine learning-based approaches, where graph vertices are mapped to a low-dimensional vector space based on local characteristics and the embedding is then optimized by reconstructing the graph from it and checking how much it deviates from the original [70, 71, 72, 73]. Such methods can be used to derive similarity measures based on the distance between two embeddings [74]; whether or not they are also applicable in lieu of normalized topological indices—in particular when it comes to their interpretability and the ability to compare structures of different size—requires a carefully set-up comparison and remains a topic for future work [75].
@@
Supplementary Material
See the supplementary material for definitions of topological indices and connections between them; details on normalization of topological indices through their cumulative distribution functions; details on characterization of the $(\mathrm{ABC}^{*},H_{2}^{*})$ space; methods of mapping RNA secondary structure to trees; methods of coarse-graining trees; extending the phase space of topological indices to graphs; and details on the $(W^{*},R^{*})$ space.
Acknowledgments
D.V. and A.B. acknowledge support by Slovenian Research Agency (ARRS) under contracts no. P1-0055 and no. J1-60002. J.M. and L.T. acknowledge support by ICSC—Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union—NextGenerationEU. J.M. acknowledges support by the European Union under NextGenerationEU grant PRIN 2022 PNRR Prot P2022TX4FE. J.M. and L.T. also thank the University of Trento for the strategic project “AIACE” that enabled this collaboration.
Author declarations
Conflict of interest
The authors declare no competing interest.
Author contributions
Domen Vaupotič: Conceptualization (equal); Formal analysis (lead); Investigation (equal); Methodology (equal); Visualization (lead); Writing – review & editing (equal). Jules Morand: Conceptualization (equal); Investigation (equal); Methodology (equal); Writing – review & editing (equal). Luca Tubiana: Conceptualization (equal); Investigation (equal); Methodology (equal); Writing – review & editing (equal). Anže Božič: Conceptualization (equal); Investigation (equal); Methodology (equal); Supervision (lead); Writing – original draft (lead); Writing – review & editing (equal).
Data and code availability
Appendix: Extended Methods
@@ Acknowledgments
The distribution parameters $p$ are first estimated with maximum likelihood estimation on sets of $10^{4}$ random trees at various $N$ ranging from $N=10$ to $N=300$ (Fig. S3 in SI). Sampling the values of topological indices on $10^{4}$ random trees is sufficient to obtain a good approximation to the distributions of the indices over the space of all (non-isomorphic) trees (Fig. S4 in SI). Afterwards, we fit the dependency of each distribution parameter $p$ on tree size $N$ with a power-law model of the form $p=\alpha N^{\beta}$ (Fig. S3). The power-law estimates of the distribution parameters are then used to obtain the cumulative distribution function of a given topological index for a tree of arbitrary size $N$.
Tree transformations
We employ four different tree transformations to study where the transformed trees lie in the phase space of normalized topological indices. Namely, we use:
  • Random leaf remove: A randomly selected leaf (degree 1 node) is removed;
  • Random leaf move: A randomly selected leaf (degree 1 node) is removed and a new leaf is added to a randomly selected node in a tree;
  • Prüfer shuffle: The Prüfer sequence of a tree is randomly shuffled, leading to a tree with a same distribution of node degrees $d(i)$ but a different topology (see, e.g., Refs. [79, 21]);
  • Random subtree swap: A randomly selected edge is removed, resulting in two disconnected trees; a new edge is then added between two randomly selected nodes of the two trees to form a new tree.